OpenStudy (anonymous):

Chris Johnson of the Tennessee Titans had 1,626 yards rushing through the first 13 games of the 2009 season. The record for rushing yards in a season is 2,105 by Eric wingspanerson. How many yards did Johnson need to average during the final 3 games in order to surpass wingspanerson's record?

OpenStudy (jdoe0001):

hmmm how many are in total in a football season anyway? I mean, Johnson did 1626 in 13 games is that the total, are there more games in order to surpass wingspaners <-- somehow I think that' snot his name anyway it'd be the difference between the quantities +1, divided by 3 that'd be the average

OpenStudy (texaschic101):

2105 - 1626 = 479 (he needs 479 more points) the average for 3 games can be found by dividing by 3... 479/3 = 159.66 rounds to 160 check.. 2105 = 1626 + 3(160) 2105 = 1626 + 480 2105 = 2106 so basically, he would have to average at least 160 per game for the last 3 games to surpass the record